# A Comprehensive Survey of Small Language Models in the Era of Large Language Models: Techniques, Enhancements, Applications, Collaboration with LLMs, and Trustworthiness

### v1

@@ 4.2.2. Runtime Efficiency Optimization
Runtime efficiency involves minimizing inference latency by reducing computing workload, decreasing the transfer of weights between GPU and CPU memory in MoE, and distributing LLM weights across multiple trusted devices.
Reducing Computing Workload. The goal of decreasing computing workload aligns with enhancing memory efficiency through methods such as quantization and KV cache compression, as mentioned in the previous section. Less model weight precision or fewer model weights naturally reduce latency. For example, quantization on an RP 3B device can achieve an inference time of 1.06 seconds per sample at 8-bit, compared to 1.66 seconds per sample at 32-bit (Rahman et al., 2023). Additionally, DMC demonstrates that $4\times$ compression can result in a $3.4\times$ acceleration on Llama 2 7B, and on an A100, the generation of 1K tokens with 3K tokens of context (Nawrot et al., 2024).
@@
Beyond quantization and KV cache compression, model collaboration, deploying smaller SLMs on devices with cloud-based LLM support, enhances runtime efficiency. Large LLMs will increase latency when directly deployed via mobile engines like llama.cpp due to a large number of computing operations. LLMCad (Xu et al., 2023d) addresses this by using a real-time, memory-resident SLM for simple tokens such as determiners and punctuation. This SLM generates tokens, while a larger cloud-based LLM verifies and corrects them, speeding up the process. LLMCad boosts token generation by up to $9.3\times$, for instance, by utilizing Llama 2 7B as the memory-resident LLM and Llama 2 13B as the cloud-based LLM, reducing latency from 16.2 seconds to 3.9 seconds on the Xiaomi Pro during TruthfulQA (Lin et al., 2022a) tasks.
@@
Reducing MoE Switching Time. To reduce latency in MoE architectures caused by frequently switching experts in limited device memory, EdgeMoE (Yi et al., 2023) enhances runtime efficiency by preemptively predicting which expert will be needed, based on the observed long-tail distribution of unbalanced expert activations. It utilizes a statistical model, built offline, to estimate expert activation probabilities in transformer layers from previous activations. During inference, EdgeMoE preemptively loads the most likely needed expert, accelerating inference by $1.11\times$ to $2.78\times$ and significantly reducing latency. For instance, in a switch transformer model with 8 experts, latency drops from approximately 0.7s to 0.3s, outperforming the baseline method that preloads experts based on hit ratios.
@@
Reducing Latency in Distributed SLMs. Distributing an SLM across smaller devices reduces the need for extensive model compression while preserving accuracy. However, this approach faces challenges that incur latency such as managing diverse device capabilities, handling data dependencies between model segments, and adapting to dynamic resource availability. To address these issues, LinguaLinked (Zhao et al., 2023a) addresses these issues by optimizing model assignment to match device capabilities and minimize data transmission, implementing runtime load balancing to redistribute tasks and prevent bottlenecks, and enhancing communication for efficient data exchange between segments. With multi-threading, the system improves, achieving acceleration rates between $1.73\times$ and $2.65\times$ for both quantized and full-precision models.
@@
Insights: We draw several key insights from the deployment of SLMs: Model size remains a bottleneck for both memory and runtime efficiency. A common solution is model quantization, which reduces model precision to save memory and lessen computing workload, thereby boosting inference speed (Yu et al., 2024c; Zhao et al., 2024c; Lin et al., 2024c; Murthy et al., 2024; Rahman et al., 2023; Liu et al., 2024h). Similarly, KV cache compression also helps achieve these efficiency gains (Kang et al., 2024; Nawrot et al., 2024; Li et al., 2024f; Yin et al., 2024). Mixture of Experts (MoE) is commonly used in SLMs to enhance performance using the same computing resources, but it results in increased memory usage. To address this, only activated experts are loaded into the memory buffer while the majority are stored cold on disk. However, the cost of switching can slow down inference. Designing a preemptive expert pre-load strategy could therefore accelerate the inference (Yi et al., 2023). Model collaboration between local SLMs and cloud-based LLMs enhances both memory and runtime efficiency by using smaller models on local devices, which are then verified by cloud LLMs to ensure performance is maintained. Using SLMs locally reduces memory usage and shortens the inference time from the local model. However, internet latency and delays in cloud LLM inference can still introduce latency. Verifying SLM outputs every NN tokens using LLMs can effectively mitigate this latency (Xu et al., 2023d). One deployment approach involves deploying SLMs/LLMs across multiple trusted local devices to maintain original performance while only loading a fraction of the model weights. However, this method can incur latency due to varying device capabilities and resource scheduling challenges. To address these issues, optimizing model assignment to align with device capabilities and minimizing data transmission are effective strategies (Zhao et al., 2023a).
5. Generic, Task-specific, and Domain-Specific Small Language Models
@@
This section investigates small language models (with fewer than 7 billion parameters) in both general and specific domains. It details the methods of obtaining these small language models, the datasets, and the evaluation tasks, exploring the techniques for acquiring SLMs through compression, fine-tuning, or training from scratch. Additionally, we summarize the representative small language models, as detailed in Table 8 and 11.
@@
Table 8. High-level Overview and Training Details of Generic-domain Small Language Models. #Params means Parameter amounts.
@@
Llama 3.2  [^2] https://huggingface.co/collections/meta-llama/llama-32-66f448ffc8c32f949b04c8cf 1B; 3B 2024.9 Pre-train Generic no release (9T tokens) GQA, SiLU, Multilingual Text and code, Shared embedding, Pruning, Distillation, SFT, RLHF, RS, DPO
Qwen 1 (Bai et al., 2023) 1.8B; 7B; 14B; 72B 2023.12 Pre-train Generic no release MHA; RoPE; SwiGLU; RMSNorm
Qwen 1.5 (Bai et al., 2023) 0.5B; 1.8B; 4B; 7B; 14B; 32B; 72B 2024.2 Pre-train Generic no release MHA; RoPE; SwiGLU; RMSNorm; Multilingual support
Qwen 2 (Yang et al., 2024e) 0.5B;1.5B; 7B; 57B; 72B 2024.6 Pre-train Generic no release GQA; RoPE; SwiGLU; RMSNorm; Multilingual support
Qwen 2.5 (Yang et al., 2024e) 0.5B; 1.5B; 3B; 7B; 14B; 32B; 72B 2024.9 Pre-train Generic no release GQA; RoPE; SwiGLU; RMSNorm; Multilingual support; Larger corpus
@@
Gemma 2 (Team et al., 2024b) 2B; 9B; 27B 2024.7 Pre-train Generic Unknown GQA; RoPE; $\text{GELU}_{\text{tanh}}$; Alternating Local and Global Attention; Logit Soft-Capping; RMSNorm for Pre and Post-Normalization
@@
Rene  [^3] https://cartesia.ai/blog/on-device 1.3B 2024.5 Pre-train Generic Dolma-1.7 (Soldaini et al., 2024) Mamba-2 layers, sliding-window attention (SWA)
@@
TinyLlama (Zhang et al., 2024i) 1B 2024.1 Pre-train Generic SlimPajama (Soboleva et al., 2023) and StarCoder (Li et al., 2023a) GQA, SiLU, Fully Sharded Data Parallel (FSDP), Flash Attention (Dao, [n. d.]), xFormers (Lefaudeux et al., 2022)
@@
Phi-3 (Abdin et al., 2024) 3.8B; 7B; 14B 2024.4 Pre-train Generic Scaled-up dataset from phi-2 MHA, SiLU, RoPE, FlashAttention, Deep ZeRO Stage 2
Phi-3.5 (Abdin et al., 2024) 3.8B; 4.2B; 6.6B 2024.4 Pre-train Generic more multilingual and long-text data Multilingual; Vision; MHA, SiLU, RoPE, FlashAttention, Deep ZeRO Stage 2
OpenELM (Mehta et al., 2024) 270M; 450M; 1.1B; 3B 2024.4 Pre-train Generic RefinedWeb (Penedo et al., 2023), deduplicated PILE (Gao et al., 2020), a subset of RedPajama (Computer, 2023), and a subset of Dolma v1.6 (Soldaini et al., 2024) No biases in FC layers; Pre-norm: RMSNorm; Pos encoding: RoPE; Attention: GQA; FFN: SwiGLU; Tokenizer: LLaMA-style
MobiLlama (Thawakar et al., 2024) 0.5B; 0.8B 2024.2 Pre-train Generic LLM360 Amber (Arxiv, Book, C4, Refined-Web, StarCoder, StackExchange, and Wikipedia) GQA; SwiGLU; Parameter-sharing
MobileLLM (Liu et al., 2024h) 125M; 350M 2024.2 Pre-train Generic Unclear (1T tokens) SwiGLU FFN, deep and thin architectures, embedding sharing, and grouped query attention
StableLM (Tow et al., [n. d.]) 3B; 7B 2023.4 Pre-train Generic RefinedWeb (Penedo et al., 2023), subsets of the Pile (Gao et al., 2020), RedPajama (Computer, 2023) and the Stack (Kocetkov et al., 2022), OpenWebText (Gokaslan et al., 2019), OpenWebMath (Paster et al., 2024), and parts of CulturaX (Nguyen et al., 2024) MHA; SiLU; Fine-tuning; DPO; Self-knowledge; RoPE; LayerNorm; no Biases
@@
Cerebras-GPT (Dey et al., 2023)  [^1] M - 13B 2023.4 Pre-train Generic Pile (Gao et al., 2020) MHA; GELU; Maximal Update Parameterization
Pythia (Biderman et al., 2023) 70M - 12B 2023.4 Pre-train Generic Pile (Gao et al., 2020) MHA; GELU; Flash Attention (Dao et al., 2022); RoPE (Su et al., 2024); ZeRO (Rajbhandari et al., 2020)
BLOOM, BLOOMZ (Le Scao et al., 2023) 560M; 1.1B; 1.7B; 3B; 7.1B; 176B 2022.11 Pre-train Generic ROOTS (Laurençon et al., 2022) and 13 programming languages MHA; $\text{GELU}_{\text{tanh}}$; ALiBi Positional Embedding (Press et al., 2022), Embedding LayerNorm (Dettmers et al., 2022)
Galactica (Taylor et al., 2022) 125M; 1.3B; 6.7B; 30B; 120B 2022.11 Pre-train Scientific Open-access scientific materials (106B tokens) but not released MHA; GeLU; Learned Positional Embeddings
@@
XGLM (Lin et al., 2022b) 1.7B; 2.9B; 7.5B 2021.12 Pre-train Generic CC100-XL -
@@
Megatron-gpt2 (Shoeybi et al., 2019) 355M; 2.5B; 8.3B 2019.9 Pre-train Generic Wikipedia (Devlin et al., 2019), CC-Stories (Trinh and Le, 2018), RealNews (Zellers et al., 2019b), OpenWebtext -
MINITRON (Muralidharan et al., 2024) 4B; 8B; 15B 2024.7 Pruning and Distillation Generic - LR Scheduler: WSD; Pruning
@@
Orca (Mukherjee et al., 2023) 13B 2023.6 Distillation FLAN-v2 (Longpre et al., 2023a) Teachers are ChatGPT, GPT4, Explanation tuning; Progressive Learning
Dolly-v2 (Conover et al., 2023) 3B; 7B; 12B 2023.4 Instruction tuning Generic Databricks-dolly-15k (Conover et al., 2023) -
LaMini-LM (Kim and Rush, 2016) 61M-7B 2023.4 Distillation Generic LaMini instruction dataset -
@@
FlanT5 (Chung et al., 2024) 80M; 250M; 780M; 3B 2022.10 Instruction Tuning Generic Muffin, T0-SF, SNI and CoT Base model is T5
T5 (Raffel et al., 2020) 60M; 220M; 770M; 3B; 11B 2019.9 Pre-train Generic C4 (Raffel et al., 2020) Encoder-decoder; Multilingual; No bias in LayerNorm; LayerNorm external to residual path

### v2

@@ 4.2.2. Runtime Efficiency Optimization
The goal of decreasing computing workload aligns with enhancing memory efficiency through methods such as quantization, as mentioned in the previous section. Decreasing model weight precision or reducing the number of weights naturally lowers latency. Other runtime efficiency techniques of minimizing inference latency involve, reducing prefill latency, early exits, large and small model collaboration, decreasing switching time in MoE, and reducing latency in distributed SLMs.
Reducing prefill latency. mllm-NPU (Xu et al., 2024d) is the first LLM inference system that leverages on-device NPU (neural processing units) to reduce prefill latency and energy consumption. It incorporates a chunk-sharing graph, shadow outlier execution, and out-of-order subgraph execution to enhance NPU offloading efficiency. Experiments have shown mllm-NPU’s superior performance benefits, including up to 43.6× speedup and 59.5× energy savings.
Dynamic Early Exits A decoupled runtime saving technique is dynamic early exits. Originating from BranchyNet (Teerapittayanon et al., 2016), which introduces exit branches after specific convolution layers in the CNN model, this concept has been adapted for PLMs as Transformer layer-wise early exiting (Xin et al., 2021). Early exiting enables dynamic acceleration during inference and reduces temporal overhead by allowing exit without passing through all model layers. To address the inconsistency issue arising from exiting at different layers, COST-EFF (Shen et al., 2022) distills a multi-exit model from the original PLM.
@@
Large and Small Model Collaboration Model collaboration, deploying SLMs on devices with cloud-based LLM support, enhances runtime efficiency. LLMs will increase latency when directly deployed via mobile engines like llama.cpp due to a large number of computing operations. LLMCad (Xu et al., 2023d) addresses this by using a real-time, memory-resident SLM for simple tokens such as determiners and punctuation. The SLM generates tokens, while a cloud-based LLM verifies and corrects them, speeding up the process. LLMCad enhances token generation up to 9.3×9.3$\times$9.3 × by pairing the memory-resident SLM, Llama 2 7B, with the cloud-based LLM, Llama 2 13B, cutting latency from 16.2 to 3.9 seconds on Xiaomi Pro for TruthfulQA tasks (Lin et al., 2022a).
@@
Reducing MoE Switching Time. To reduce latency in MoE architectures caused by frequently switching experts in limited device memory, EdgeMoE (Yi et al., 2023) enhances runtime efficiency by preemptively predicting which expert will be needed, based on the observed long-tail distribution of unbalanced expert activations. It utilizes a statistical model, built offline, to estimate expert activation probabilities in transformer layers from previous activations. During inference, EdgeMoE preemptively loads the most likely needed expert, accelerating inference by $1\times1$ to 2.78×2.78$\times$2.78 × and significantly reducing latency. For instance, in a switch transformer model with 8 experts, latency drops from approximately 0.7s to 0.3s, outperforming the baseline method that preloads experts based on hit ratios.
@@
Reducing Latency in Distributed SLMs. Distributing an SLM across smaller devices reduces the need for extensive model compression while preserving accuracy. However, this approach faces challenges that incur latency such as managing diverse device capabilities, handling data dependencies between model segments, and adapting to dynamic resource availability. To address these issues, LinguaLinked (Zhao et al., 2023a) addresses these issues by optimizing model assignment to match device capabilities and minimize data transmission, implementing runtime load balancing to redistribute tasks and prevent bottlenecks, and enhancing communication for efficient data exchange between segments. With multi-threading, the system improves, achieving acceleration rates between 1.73×1.73$\times$1.73 × and 2.65×2.65$\times$2.65 × for both quantized and full-precision models.
@@
Insights: We draw several key insights from the deployment of SLMs: Model size remains a bottleneck for both memory and runtime efficiency. A common solution is model quantization, which reduces model precision to save memory and lessen computing workload, thereby boosting inference speed (Yu et al., 2024c; Zhao et al., 2024c; Lin et al., 2024c; Murthy et al., 2024; Rahman et al., 2023; Liu et al., 2024h). Similarly, KV cache compression also helps achieve these efficiency gains (Kang et al., 2024; Nawrot et al., 2024; Li et al., 2024g; Yin et al., 2024). Mixture of Experts (MoE) is commonly used in SLMs to enhance performance using the same computing resources, but it results in increased memory usage. To address this, only activated experts are loaded into the memory buffer while the majority are stored cold on disk. However, the cost of switching can slow down inference. Designing a preemptive expert pre-load strategy could therefore accelerate the inference (Yi et al., 2023). Model collaboration between local SLMs and cloud-based LLMs enhances both memory and runtime efficiency by using smaller models on local devices, which are then verified by cloud LLMs to ensure performance is maintained. Using SLMs locally reduces memory usage and shortens the inference time from the local model. However, internet latency and delays in cloud LLM inference can still introduce latency. Verifying SLM outputs every $N$ tokens using LLMs can effectively mitigate this latency (Xu et al., 2023d). One deployment approach involves deploying SLMs/LLMs across multiple trusted local devices to maintain original performance while only loading a fraction of the model weights. However, this method can incur latency due to varying device capabilities and resource scheduling challenges. To address these issues, optimizing model assignment to align with device capabilities and minimizing data transmission are effective strategies (Zhao et al., 2023a).
5. Generic and Domain-Specific Small Language Models
@@
This section investigates small language models (with fewer than 7 billion parameters) in both general and specific domains. It details the methods of obtaining these small language models, the datasets, and the evaluation tasks, exploring the techniques for acquiring SLMs through compression, fine-tuning, or training from scratch. Additionally, we summarize the representative small language models, as detailed in Table 7 and 10.
@@
Table 7. High-level Overview and Training Details of Generic-domain Small Language Models. #Params means Parameter amounts. ”¿” indicates parameters larger than 7B.
@@
PhoneLM (Yi et al., 2024) 0.5B; 1.5B 2024.11 Pre-train Generic DCLM-baseline (Li et al., 2024b), StarCoderData (Li et al., 2023a); OpenWebMath (Paster et al., 2024), Dolma-algebraic and Dolma-arXiv (Soldaini et al., 2024) RoPE, MHA, Gated FFN, RMSNorm, ReLU, FSDP, Flash Attention 2, ZeRO
Llama 3.2 (AI, 2024) 1B; 3B 2024.9 Pre-train Generic no release (9T tokens) GQA, SiLU, Multilingual Text and code, Shared embedding, Pruning, Distillation, SFT, RLHF, RS, DPO
Qwen 1 (Bai et al., 2023) 1.8B; 7B; ¿ 2023.12 Pre-train Generic no release MHA; RoPE; SwiGLU; RMSNorm
Qwen 1.5 (Bai et al., 2023) 0.5B; 1.8B; 4B; 7B; ¿ 2024.2 Pre-train Generic no release MHA; RoPE; SwiGLU; RMSNorm; Multilingual support
Qwen 2 (Yang et al., 2024h) 0.5B;1.5B; 7B; ¿ 2024.6 Pre-train Generic no release GQA; RoPE; SwiGLU; RMSNorm; Multilingual support
Qwen 2.5 (Yang et al., 2024h) 0.5B; 1.5B; 3B; 7B; > 2024.9 Pre-train Generic no release GQA; RoPE; SwiGLU; RMSNorm; Multilingual support; Larger corpus
@@
Gemma 2 (Team et al., 2024b) 2B; ¿ 2024.7 Pre-train Generic Unknown GQA; RoPE; $\text{GELU}_{\text{tanh}}$; Alternating Local and Global Attention; Logit Soft-Capping; RMSNorm for Pre and Post-Normalization
SmolLM (Allal et al., 2024) 135M; 360M; 1.7B 2024.7 Pre-train Generic smollm-corpus (Ben Allal et al., 2024) GQA, trapezoidal LR scheduler
@@
Rene (Goel, 2024) 1.3B 2024.5 Pre-train Generic Dolma-1.7 (Soldaini et al., 2024) Mamba-2 layers, sliding-window attention (SWA)
@@
CT-LLM (Du et al., 2024) 2B 2024.4 Pre-train Generic MAP-CC Chinese, MHA, RoPE, SwiGLU, RMSNorm
@@
TinyLlama (Zhang et al., 2024j) 1B 2024.1 Pre-train Generic SlimPajama (Soboleva et al., 2023) and StarCoder (Li et al., 2023a) GQA, SiLU, FSDP, Flash Attention (Dao, [n. d.]), xFormers (Lefaudeux et al., 2022)
@@
Phi-3 (Abdin et al., 2024) 3.8B; 7B; ¿ 2024.4 Pre-train Generic Scaled-up dataset from phi-2 MHA, SiLU, RoPE, FlashAttention, Deep ZeRO Stage 2
Phi-3.5 (Abdin et al., 2024) 3.8B; 4.2B; 6.6B 2024.4 Pre-train Generic more multilingual and long-text data Multilingual; Vision; MHA, SiLU, RoPE, FlashAttention, ZeRO 2
OpenELM (Mehta et al., 2024) 270M; 450M; 1.1B; 3B 2024.4 Pre-train Generic RefinedWeb (Penedo et al., 2023), deduplicated PILE (Gao et al., 2020), partial RedPajama (Computer, 2023), partial Dolma v1.6 (Soldaini et al., 2024) No biases in FC layers; Pre-norm: RMSNorm; Pos encoding: RoPE; Attention: GQA; FFN: SwiGLU; Tokenizer: LLaMA-style
MobiLlama (Thawakar et al., 2024) 0.5B; 0.8B 2024.2 Pre-train Generic LLM360 Amber  [^3] LLM360 Amber includes Arxiv, Book, C4, Refined-Web, StarCoder, StackExchange, and Wikipedia GQA; SwiGLU; Parameter-sharing
MobileLLM (Liu et al., 2024h) 125M; 350M 2024.2 Pre-train Generic Unknown (1T tokens) SwiGLU FFN, deep and thin architectures, embedding sharing, and grouped query attention
StableLM (Tow et al., [n. d.]) 3B; 7B 2023.4 Pre-train Generic RefinedWeb (Penedo et al., 2023), RedPajama (Computer, 2023), the Stack (Kocetkov et al., 2022), OpenWebText (Gokaslan et al., 2019), OpenWebMath (Paster et al., 2024), and partial CulturaX (Nguyen et al., 2024) MHA; SiLU; Fine-tuning; DPO; Self-knowledge; RoPE; LayerNorm; no Biases
@@
Cerebras-GPT (Dey et al., 2023)  [^1] M; 256M; 590M; 1.3B; 2.7B; 6.7B; ¿ 2023.4 Pre-train Generic Pile (Gao et al., 2020) MHA; GELU; Maximal Update Parameterization
Pythia (Biderman et al., 2023) 14M;70M;160M;410M; 1B;1.4B;2.8B;6.9B;¿ 2023.4 Pre-train Generic Pile (Gao et al., 2020) MHA; GELU; Flash Attention (Dao et al., 2022); RoPE (Su et al., 2024); ZeRO (Rajbhandari et al., 2020)
BLOOM, BLOOMZ (Le Scao et al., 2023) 560M; 1.1B; 1.7B; 3B; 7.1B; ¿ 2022.11 Pre-train Generic ROOTS (Laurençon et al., 2022) and 13 programming languages MHA; $\text{GELU}_{\text{tanh}}$; ALiBi Positional Embedding (Press et al., 2022), Embedding LayerNorm (Dettmers et al., 2022)
Galactica (Taylor et al., 2022) 125M; 1.3B; 6.7B; ¿ 2022.11 Pre-train Scientific Open-access scientific materials (106B tokens) but not released MHA; GeLU; Learned Positional Embeddings
@@
XGLM (Lin et al., 2022b) 1.7B; 2.9B; ¿ 2021.12 Pre-train Generic CC100-XL -
@@
Megatron-gpt2 (Shoeybi et al., 2019) 355M; 2.5B; ¿ 2019.9 Pre-train Generic Wikipedia (Devlin et al., 2019), CC-Stories (Trinh and Le, 2018), RealNews (Zellers et al., 2019b), OpenWebtext -
MINITRON (Muralidharan et al., 2024) 4B; ¿ 2024.7 Distillation; Pruning Generic 8T tokens in Nemotron-4 (Parmar et al., 2024) LR WSD Scheduler
@@
Orca (Mukherjee et al., 2023) 13B 2023.6 Distillation FLAN-v2 (Longpre et al., 2023a) From ChatGPT and GPT4, Explanation tuning; Progressive Learning
MINIMA (Zhang et al., 2023b) 3B 2023.11 Distillation Generic Pile (Gao et al., 2020), Wudao (Yuan et al., 2021), GitHub (Computer, 2023) From Llama-2-7B, Zero2, Flash Attention, Optimal teacher size
Dolly-v2 (Conover et al., 2023) 3B; 7B; ¿ 2023.4 Instruction tuning Generic Databricks-dolly-15k (Conover et al., 2023) from pythia
LaMini-LM (Kim and Rush, 2016) 61M-7B 2023.4 Distillation Generic LaMini instruction dataset a collection of SLMs distilled from ChatGPT-generated 2.58M instructions.